### v2

# ShapeICP: Iterative Category-Level Object Pose and Shape Estimation from Depth

Abstract
Category-level object pose and shape estimation from a single depth image has recently drawn research attention due to its potential utility for tasks such as robotics manipulation. The task is particularly challenging because the three unknowns, object pose, object shape, and model-to-measurement correspondences, are compounded together, but only a single view of depth measurements is provided. Most of the prior work heavily relies on data-driven approaches to obtain solutions to at least one of the unknowns, and typically two, running with the risk of failing to generalize to unseen domains. The shape representations used in the prior work also mainly focus on point cloud and signed distance field (SDF). In stark contrast to the prior work, we approach the problem using an iterative estimation method that does not require learning from pose-annotated data. In addition, we adopt a novel mesh-based object active shape model that the previous literature has not explored. Our algorithm, ShapeICP, is based on the iterative closest point (ICP) algorithm but is equipped with additional features for the category-level pose and shape estimation task. Although not using pose-annotated data, ShapeICP surpasses many data-driven approaches that rely on pose data for training, opening up a new solution space for researchers to consider.
I Introduction
We investigate the task of estimating the pose and shape of an object given a single depth image of it. The object instance is only known up to its semantic category. Its exact geometry (i.e. CAD model) is unavailable to us. This problem is common in many scenarios, from warehouse robots to home robots. For example, home robots may require estimating the poses and shapes of tabletop objects to manipulate them. Because of the significance of this task, it has drawn substantial research attention in recent years.
The category-level object pose and shape estimation problem is typically set up such that the instance segmentation and semantic classification of the object are performed on the accompanying RGB image with an off-the-shelf method [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. However, even after factoring out the object segmentation problem, inferring the object pose and shape from the segmented depth image remains challenging due to the three entangled unknowns: object pose, object shape, and model-to-measurement correspondences. It requires us to know two of the three unknowns to reduce the problem to an easily solvable one. For instance, if we know the object shape and how the measurements correspond to the points on the shape, it is reduced to the problem of aligning two point sets, which the Umeyama algorithm [13] can solve. If the object pose and the correspondences of measurements to a shape template are known, one can transform and deform the template to the measurements to obtain the shape. Researchers often recognize the correspondences also as unknowns and deform the template [14] using Chamfer loss, where the closest points are the correspondences [15, 16, 14]. However, with the three variables all unknown, the problem is extremely difficult.
Most prior methods exploit neural networks to learn a mapping from inputs such as the RGB-D image and a categorical shape prior to one or two of the three unknowns. Although these learning methods greatly factor away the complexities in the problem by leveraging data and have achieved excellent performance on benchmarks, they face potential failures caused by domain changes and the tedious data curation process required for training when they are deployed to a new environment. The ground truth pose of an object is especially painstaking to annotate since so far there has been no scalable and easily accessible approach to measure the pose of an object. On the other hand, virtually all the prior methods focus on either point cloud or signed distance field (SDF) as the shape representation. Point clouds are flexible but do not have the notion of surface, while SDF entails unfavorable computational complexity, which is often scaled cubically with the resolution.
In this work, we take a radically different approach. First, in contrast to prior methods, we adopt a novel mesh-based active shape model (ASM) [17] as the object shape representation. Meshes incorporate the notion of surface, being superior to point clouds. They are also inherently a geometric representation, unlike an SDF or occupancy grid where the geometry is represented implicitly and some processing, such as finding the zero level set (SDF) or marching cubes [18] (occupancy grid), is required to extract the geometry. Using meshes in category-level object pose and shape estimation is also a relatively untapped terrain in the literature; it is worth exploring more. Second, although methods such as self-supervised learning and sim-to-real transfer can help circumvent the annotation difficulty, we seek a different avenue by devising an iterative estimation algorithm based on the iterative closest point (ICP) algorithm [19] and many other classical estimation techniques without training any module on the pose-annotated data. The method has the option of using an image-based shape classification network for shape initialization to boost estimation performance. We consider that in practice, the ground-truth shape of an object is more accessible than its ground-truth pose since object scanning and depth fusion techniques [20] are mature.
We evaluate our method on the NOCS REAL dataset [21] and find it better or comparable to many learning-based methods, even though our method is at the disadvantage that no pose-annotated data are used. The evaluation of our method also reveals what can already be accomplished without learning and why learning may be needed in some places. These insights are traditionally difficult to draw from the evaluation of black-box style learning-based methods.
II Related Work
We will review the related work in terms of which of the three unknowns, object pose, object shape, and model-to-measurement correspondences, are predicted by neural networks. We will also discuss shape representations, particularly ASM, used in the literature.
II-A Category-Level Pose and Shape Estimation
Predicting shape and correspondences. The seminal work NOCS [21] belongs to this category. The predicted NOCS map encodes the depth profile (i.e. shape) of the observed object in its canonical pose. The NOCS map also corresponds pixel-to-pixel to the measured depth image. Another work is [11] where the authors first use Mesh R-CNN [15] to predict a metric-scale mesh of the object shape from the image and then predict the NOCS map to solve for the pose. A notable line of work [1, 2, 3] attempts to predict the deformation applied to a categorical shape prior to obtain the object shape and a correspondence matrix to associate the model (i.e. the deformed prior) points to the measured depth points. For all these methods, the final step is the Umeyama algorithm [13] to solve for the object pose.
Predicting shape and pose. CASS [4] first learns a point cloud based shape embedding space. During inference, the shape latent code is predicted given the RGB-D image patch. The predicted latent code is both decoded to the shape and fed to the module for pose prediction. CenterSnap [22] and ShAPO [23] build a YOLO-style [24] one-shot estimation pipeline. iCaps [5] selects the rotation by comparing the depth image features to the features in a pre-computed code book and predicts the DeepSDF [25] latent vector as the shape representation. Other methods that directly regress shape and pose to various extents include [6, 7, 26].
Predicting pose. This category of methods bypasses shape estimation and estimates only the pose. FS-Net [27] extracts rotation-aware features through a scale and shift-invariant 3D graph convolution network. It decouples the rotation into two perpendicular vectors for prediction. SAR-Net [8] trains neural networks to transform a categorical template to the same rotation as the observed point cloud and complete the observed point cloud by mirroring about the symmetry plane. These network outputs are used to compute the pose.
Predicting correspondences. [28] assumes known correspondences provided by neural keypoint matching and solves an alignment problem between the model keypoints and the measurement keypoints. Instead of only estimating the $\mathrm{SIM}(3)$ pose as in Umeyama [13], the method additionally estimates the parameters for an active shape model (ASM) [17] of the object. The ASM is linear and can be treated similarly to translation. The method also features a graph-theoretic approach and an outlier-robust solver to handle false correspondences. [29] is an earlier work that essentially shares the same problem formulation as [28] but with a different solver that does not consider the outlier issue.
Optimization-based. Neural Analysis-by-Synthesis [9] leverages neural rendering to render an image from the object shape code and pose. The rendered image is compared with the actual image to optimize the shape code and pose. DirectShape [10] sets up the optimization losses by silhouette matching and left-right stereo photo consistency using an SDF-based ASM. Strong regularization, such as the ground plane constraint, is engaged to alleviate the optimization difficulty. ELLIPSDF [12] employs a bi-level shape representation to coarsely fit the object segmentation for initialization and then finely optimize the pose and shape.
As we have seen, previous methods commonly adopt data-driven solutions. Pose-annotated data are exploited in almost all the prior work, either explicitly or implicitly. Even the optimization-based methods may actually require extra information such as multiple views [12] or application-specific constraints [10]. Our method works completely without pose-annotated data. The addition of an optional shape classification network merely requires shape-annotated data, which also falls into the rare category of predicting only the shape, a direction not well explored by the prior work.
II-B Shape Representation – Active Shape Model
The active shape model (ASM) [17] has been employed in the category-level object pose and shape estimation task [10, 29, 28, 6], and also in the 3D object shape reconstruction task [30]. The ASM has the advantage of being a simple linear model that is more efficient to back-propagate through during optimization compared to a neural network. The object ASM is primarily point-based [29, 28, 6] and SDF-based [10, 30]. It is interesting to explore the mesh representation further.
III Method
III-A Overview
Our task is to estimate the $\mathrm{SIM}(3)$ pose and shape of an object which is only known up to its semantic category given a single depth image of it. We assume that instance segmentation of the object and its class label are provided by an off-the-shelf method such as Mask RCNN [31] as typically used in the prior work [9, 10, 1, 2, 3, 4, 6, 5, 7, 8]. We adopt a mesh-based object active shape model (ASM) [17] for its flexibility and surface representability. After obtaining the ASM offline (Section III-B), we estimate the pose and shape by transforming and fitting the shape model to the observed depth points. The optimization is realized by an augmented iterative closest point (ICP) algorithm which, in addition to the existing pose estimation step, has a shape deformation step for shape estimation (Section III-C). We thus name our algorithm ShapeICP. However, a significant challenge is that ICP is a local solver, and so is the base form of ShapeICP. To overcome this, ShapeICP uses multi-hypothesis tracking for rotation estimation, expectation maximization (EM) for robust correspondence handling, and shape classification for shape initialization (Section III-D). The full pseudo-algorithm of ShapeICP can be found in Section III-E.
III-B Mesh-based Object Active Shape Model
Template deformation. Mesh-based ASMs have been applied to human faces [32] but not yet in the task of object pose and shape estimation. The bottleneck is that object models from a database have different numbers of vertices and different vertex connectivity even within the same category, making it difficult to draw corresponding vertices across models for PCA. To solve this challenge, we deform a template mesh to wrap around each object model and run PCA on the vertices of the deformed templates for all the object models within a category. Specifically, we denote a mesh by $\mathcal{M=(V,E,F)}$, where $\mathcal{V}=\{v_{i}\}_{i=1}^{V}$ are the $V$ vertices in the mesh, =$\mathcal{E}=\{e_{i}\}_{i=1}^{E}$ is the set of $E$ tuples of the indices of a pair of vertices connected by an edge, =$\mathcal{F}=\{f_{i}\}_{i=1}^{F}$ is the set of $F$ tuples of indices of vertices on the same face. To deform the template (source) mesh $\mathcal{M}_{s}$ to a target mesh $\mathcal{M}_{t}$, we randomly sample points on the mesh surface with $N$ points from the template mesh $p_{n}\sim\mathcal{M}_{s}$ and $M$ points from the target mesh $q_{m}\sim\mathcal{M}_{t}$ (here $N=M$). The deformation is done by minimizing a similar set of losses as in [16, 15, 33]:
$\mathcal{L}_{c}=\frac{1}{N}\sum_{n}{\min_{m}{\|p_{n}-q_{m}\|_{2}^{2}}}+\frac{1}{M}\sum_{m}{\min_{n}{\|p_{n}-q_{m}\|_{2}^{2}}}$ (1)
which is the Chamfer distance between the two sets of sampled points. To regularize the template mesh during the deformation process, a normal consistency loss is imposed on the template mesh:
$\mathcal{L}_{n}=\frac{1}{E}\sum_{e}{1-\cos{(\mathbf{n}_{f_{e+}},\mathbf{n}_{f_{e-}})}}$ (2)
where $\mathbf{n}_{f_{e+}}$ and $\mathbf{n}_{f_{e-}}$ are the normals of the two faces that share edge $e$, and $\cos$ is the cosine distance. This loss regularizes the surface normals to be smooth. To avoid overly long outlier edges in the template mesh, an edge length loss is added:
$\mathcal{L}_{e}=\frac{1}{E}\sum_{e}{\|v_{e+}-v_{e-}\|_{2}^{2}}$ (3)
where $v_{e+}$ and $v_{e-}$ are the vertices connected by the edge $e$. A Laplacian smoothing loss is also enforced on the template mesh to encourage the vertices to move along with their neighbors and potentially avoid mesh self-intersection:
$\mathcal{L}_{l}=\frac{1}{V}\sum_{i}{\|v_{i}-\frac{1}{|\mathcal{N}(i)|}\sum_{j\in\mathcal{N}(i)}{v_{j}}\|_{2}}$ (4)
where $\mathcal{N}(i)=\{j\}~{}\forall(i,j)\in\mathcal{E}$ is the set of neighboring vertices of vertex $v_{i}$. The overall objective function is $\min_{\mathcal{V}_{s}}\mathcal{L}_{c}+\lambda_{n}\mathcal{L}_{n}+\lambda_{e}\mathcal{L}_{e}+\lambda_{l}\mathcal{L}_{l}$, where $\lambda$’s are the weights for the loss terms. This optimization is carried out by stochastic gradient descent (SGD) [34] starting from a spherical template with $p_{n}$ and $q_{m}$ resampled in each step. An example of the process is visualized in Fig. 1.
Figure 1: An example of the template deformation process with the initial template on the left and the target mesh on the right side of the arrow.
PCA. Now that we have acquired a set of deformed templates that take on the model shapes across a category and have a consistent number of vertices and connectivity, we can build a feature vector for each model by concatenating the vertices of its deformed template, i.e. $[v_{1}...v_{V_{s}}]$, and then running PCA on all the feature vectors to obtain our category-level mesh-based ASM. A vertex in the final ASM with $K$ principal components is expressed as
$v_{i}=b_{0,i}+\sum_{k=1}^{K}{c_{k}b_{k,i}}$ (5)
where $b_{0,i}$ is the vertex in the mean shape corresponding to $v_{i}$, $b_{k,i}$ is the corresponding vertex in basis $k$, and $\{c_{k}\}_{k=1}^{K}$ are the weights for the bases. These weights parameterize the shape. Fig. 2 visualizes the vertices of the bases and the final shape in an example ASM. The edges and faces of the mesh-based ASM are inherited from the template mesh.
Figure 2: Visualization of the vertices of the bases and the final shape in an example ASM. The final shape on the right shows how the mean shape is displaced by the weighted bases (the scale of the bases is much smaller than the mean, so the bases on the left are enlarged for clarity).
III-C Alternating Pose and Shape Optimization
Formulation. To introduce our ShapeICP formulation, we need sample points on the ASM. Each sampled point can be expressed as
$p_{n}=\sum_{i\in\mathsf{F}(n)}{\mu_{n,i}v_{i}}~{}~{}\text{s.t.}\sum_{i\in\mathsf{F}(n)}{\mu_{n,i}}=1$ (6)
where $\mathsf{F}(n)$ is a function that maps the sample index $n$ to a face randomly selected with probability proportional to the face area, and $\mu_{n,i}$ are randomly generated interpolation weights for this sample. These interpolation weights are fixed during the later optimization. Combining with (5), $p_{n}=\sum_{i\in\mathsf{F}(n)}{\mu_{n,i}(b_{0,i}+\sum_{k=1}^{K}{c_{k}b_{k,i}})}$. We now introduce the ShapeICP formulation:
$\min_{\begin{subarray}{c}R\in\mathrm{SO}(3),t\in\mathbb{R}^{3},s\in\mathbb{R}_{++}\\ \mathbf{c}=[c_{1}...c_{K}]\in\mathbb{R}^{K}\end{subarray}}{\frac{1}{M}\sum_{m=1}^{M}{\min_{n}{\|sRp_{n}+t-q_{m}\|_{2}^{2}}}}$ (7)
where $\{q_{m}\}_{m=1}^{M}$ (overloaded notation) are now the object depth points segmented from the depth image and back-projected to 3D using the camera intrinsics, $R$, $t$, and $s$ are the object rotation, translation, and scale to be estimated, $\mathbf{c}$ which enters the objective function through $p_{n}$ is the ASM shape code also to be estimated. For each measurement $q_{m}$, (7) encourages the transformed model points to be close to the measurements, which is similar to the ICP objective function [19] but with the addition of the shape parameter $\mathbf{c}$. We also remark that formulation (7) equivalently works for a point-based ASM where $p_{n}$ is simply a point in the ASM point cloud. However, a mesh-based ASM has the notion of surface and thus is a better geometric representation. Furthermore, a mesh model enables the possibility of rendering to be used in Section III-D.
Core algorithm. We make two observations of (7). First, if the shape code $\mathbf{c}$ is given, the problem shares the same objective function as ICP. Second, if the $\mathrm{SIM}(3)$ pose is given, the problem mimics the mesh deformation problem introduced in Section III-B except that (7) is only one-sided Chamfer distance and the optimization variable is the shape code instead of the vertices. These two observations lead to an intuitive alternating minimization algorithm to solve (7):
  • Pose step: Associate each measurement $q_{m}$ to the closest transformed model point $\hat{s}\hat{R}\hat{p}_{n}+\hat{t}$ (i.e. solving $\min_{n}$ in (7)) using the last estimates $\hat{\mathbf{c}}$, $\hat{R}$, $\hat{t}$, and $\hat{s}$. Solve for the incremental $R$, $t$, and $s$ with Umeyama [13] to further align the model points with the measurements while keeping $\mathbf{c}$ fixed.
  • Shape step: Re-associate each $q_{m}$ to a model point using the latest estimates. Solve for $\mathbf{c}$ as if it is a mesh deformation problem while keeping $R$, $t$, and $s$ fixed.
After the pose step, the incremental $R$, $t$, and $s$ are accumulated onto the estimates $\hat{R}$, $\hat{t}$, and $\hat{s}$. Every time $\mathbf{c}$ is updated, $p_{n}$’s are re-sampled to ensure sample uniformity as the areas of the mesh faces change.
Shape step. In the shape step, we include the regularization losses (2) – (4). The final objective function $\mathcal{L}_{sc}$ ($sc$: shape code) in the shape step is $\min_{\mathbf{c}}{\mathcal{L}_{sc}}=\min_{\mathbf{c}}\mathcal{L}_{ps}+\lambda_{n}\mathcal{L}_{n}+\lambda_{e}\mathcal{L}_{e}+\lambda_{l}\mathcal{L}_{l}$, where $\mathcal{L}_{ps}$ ($ps$: pose shape) refers to the objective function in (7) and $\lambda$’s are the weights. These weights can be reused from Section III-B if we center and normalize the observed point cloud, and shift and normalize the transformed model points by the same amount (because the ShapeNet [35] objects in Section III-B are all centered and normalized). Since this is an iterative coordinate-descent style algorithm, we prefer small increments on $\mathbf{c}$. Only one step of gradient descent is performed on $\mathbf{c}$ during each shape step. However, we repeat the shape step several times in each iteration. The shape step is skipped in the final iterations for better empirical performance.
III-D Coping with Local Minima
Category-level object pose and shape estimation given only a single-view depth image naturally comes with large ambiguity (Fig. 3) since many configurations of the variables may fit equally well to the observed partial object point cloud. In other words, the optimization landscape exhibits many local minima, which hurt the performance of a local solver such as ShapeICP. In this section, we develop several strategies to cope with the local minima.
Figure 3: An occluded depth observation of the laptop (whose point cloud is fragmented in the right picture) causes ambiguity. The fitted model (whose vertices are visualized as the organized point set) has pose errors and collides with another object.
Expectation maximization (EM). The correspondences between the measured depth points and the model points can be treated as latent variables. Instead of hard one-to-one associations, EM allows softer associations [36], reducing the chance of being stuck at a local minimum. Inspired by [37], we integrate EM into our ShapeICP formulation. The new objective function $\mathcal{L}_{em}$ derived from EM is (details about the probability models, independence assumptions, and derivation can be found in the Appendix):
$\min_{\begin{subarray}{c}R\in\mathrm{SO}(3),t\in\mathbb{R}^{3},s\in\mathbb{R}_{++}\\ \mathbf{c}=[c_{1}...c_{K}]\in\mathbb{R}^{K}\end{subarray}}{\sum_{m=1}^{M}{\sum_{n\in\mathcal{N}_{Q}(q_{m})}{w_{mn}\|sRp_{n}+t-q_{m}\|_{2}^{2}}}}$ (8)
where
$w_{mn}=\frac{\mathsf{N}(\hat{s}\hat{R}\hat{p}_{n}+\hat{t}-q_{m};0,\sigma_{m})}{2\sigma^{2}_{m}\sum_{n\in\mathcal{N}_{Q}(q_{m})}{\mathsf{N}(\hat{s}\hat{R}\hat{p}_{n}+\hat{t}-q_{m};0,\sigma_{m})}}$ (9)
where the hatted quantities $\hat{s}$, $\hat{R}$, $\hat{t}$, and $\hat{p}_{n}$ ($\hat{p}_{n}$ as a result of $\hat{\mathbf{c}}$) are the estimates from the last step and do not participate in the current optimization, $\mathcal{N}_{Q}(q_{m})$ finds the $Q$ closest transformed model points to $q_{m}$ (i.e. the top $Q$ solutions to $\arg\min_{n}{\|sRp_{n}+t-q_{m}\|_{2}^{2}}$), and $\mathsf{N}(\cdot;0,\sigma_{m})$ is a zero-mean Gaussian distribution with isotropic covariance $\Sigma_{m}=\mathsf{diag}(\sigma_{m}^{2})$. EM yields an intuitive formulation that considers $Q$ likely correspondences weighted by their probabilities computed from the last-step estimates rather than one hard correspondence. This increases the field of view of the model-measurement association step and can help move out of local minima. Both the pose and shape steps can use (8).
Multi-hypothesis estimation. Initialization is critical in avoiding local minima. The translation can be initialized as the centroid of the observed point cloud $\bar{q}=\frac{1}{M}\sum_{m}{q_{m}}$. We find it better to begin with a small initial scale such as the average distance to the center $\frac{1}{M}\sum_{m}{q_{m}-\bar{q}}$. Without any neural network, the shape code is initialized to be the mean code for the category $\bar{\mathbf{c}}=\frac{1}{U}\sum_{u}{\tilde{\mathbf{c}}_{u}}$ where $U$ is the total number of database object models for the category. However, the rotation initial guess is not readily available. We thus track multiple rotation hypotheses in parallel and drop unpromising hypotheses quickly to save computation. Specifically, we start with the base $\mathrm{SO}(3)$ grid from [38], as also used by [26]. The grid has 2304 discrete rotations covering $\mathrm{SO}(3)$. During the course of optimization, we drop hypotheses according to the following score functions:
Objective function. The mean residual in (7) $\mathcal{S}_{r}=\frac{1}{M}\sum_{m}{r_{m}}=\frac{1}{M}\sum_{m}{\min_{n}{\|\hat{s}\hat{R}\hat{p}_{n}+\hat{t}-q_{m}\|_{2}^{2}}}$ plus the standard deviation of the residuals $\mathcal{S}_{\sigma}=\mathsf{std}(\{r_{m}\}_{m=1}^{M})$ is used. We find that adding the standard deviation improves the results, as we hope to have relatively uniform residuals over all the depth points.
Symmetry check. Man-made objects often have mirror symmetry (e.g. laptops) and rotational symmetry (e.g. bottles and cans). If the estimated pose and shape are correct, manipulating the observed point cloud with respect to the estimated plane and axis of symmetry should still result in a low residual. Let $\hat{T}=[\hat{s}\hat{R}~{}\hat{t};~{}0~{}0~{}0~{}1]$ be the $\mathrm{SIM}(3)$ pose estimate, and $\hat{p}_{h,n}=[\hat{p}_{n};1]$ and $q_{h,m}=[q_{m};1]$ be the homogeneous coordinates. The symmetry score $\mathcal{S}_{\psi,r}$ is
$\frac{1}{M}\sum_{m}{r_{\psi,m}}=\frac{1}{M}\sum_{m}{\min_{n}{\|\hat{T}\hat{p}_{h,n}-\hat{T}T_{\psi}\hat{T}^{-1}q_{h,m}\|_{2}^{2}}}$ (10)
where $T_{\psi}=[R_{\psi}~{}0;~{}0~{}0~{}0~{}1]$ is a symmetry operation such as rotation around the axis of symmetry or reflection with respect to the plane of symmetry (in which case $\det(R_{\psi})=-1$). (10) essentially transforms the observed point cloud to the object canonical pose using the current estimate $\hat{T}$, applies $R_{\psi}$, and transforms back so that the order of magnitude of $\mathcal{S}_{\psi,r}$ is similar to $\mathcal{S}_{r}$. The standard deviation score is $\mathcal{S}_{\psi,\sigma}=\mathsf{std}(\{r_{\psi,m}\}_{m=1}^{M})$ and the total symmetry score is $\mathcal{S}_{\Psi}=\frac{1}{\Psi}\sum_{\psi}{\mathcal{S}_{\psi,r}+\mathcal{S}_{\psi,\sigma}}$ for a total of $\Psi$ different symmetry operations (multiple rotation symmetry operations of different angles are possible).
Depth rendering. Taking advantage of the mesh-based representation, we render the current estimates to a depth image and compare with the observed depth image of the object (after applying the segmentation mask). The depth rendering score is
$\mathcal{S}_{dr}=\frac{1}{|\Omega(D)|}\sum_{d\in\Omega(D)}{[\mathsf{R}(\hat{T},\mathcal{M}(\hat{\mathbf{c}}),K_{c})_{d}-D_{d}]^{2}}$ (11)
where $D$ is the masked observed depth image, $d\in\Omega(D)$ indexes pixels in the image space, $\mathsf{R}$ is the renderer, $\mathcal{M}(\hat{\mathbf{c}})$ is the mesh model built from $\hat{\mathbf{c}}$, and $K_{c}$ includes the known camera intrinsics and other camera parameters (such as image size). We use the renderer from [33] and assign the same background value to the rendered image and the masked observed depth image. The depth rendering helps to identify an over-sized model (which may still give low $\mathcal{S}_{r}$ if a portion of it fits the measurements well) [39]. To avoid computation overflow, we start this depth rendering after the number of hypotheses is low.
The total score to be minimized is $\mathcal{S}_{tot}=\mathcal{S}_{r}+\mathcal{S}_{\sigma}+\lambda_{\Psi}\mathcal{S}_{\Psi}+\lambda_{dr}\mathcal{S}_{dr}$ where $\lambda$’s are the weights. When picking top hypotheses, we greedily select the next hypothesis at least some angle ($\theta$) apart from the last chosen hypothesis to alleviate duplicated or very close hypotheses.
Shape classification. To better initialize the shape code $\mathbf{c}$, we build a shape classification network whose input is a color-coded normal vector image of the segmented object and the output is the index of the closest database model $u^{*}$. We choose ResNet-50 [40] for the network. The normal vector image is computed from the observed depth image and converted from $x,y,z\in[0,1]$ (assuming a unit vector) to RGB values. Only the image patch that contains the object (i.e. the 2D bounding box) is resized to a constant size and fed to the network. We find the ground-truth closest model by computing Chamfer distance (1) between the ground-truth shape and all the models in the database (ShapeNet [35]). The final initial guess is $\tilde{\mathbf{c}}_{u^{*}}$ for the shape code. Note that if a model in the database is not matched to any ground-truth shape, it does not appear as a class in the network output. Therefore, the number of classes is smaller than or equal to the number of ground-truth shapes.
III-E Summary
The full estimation algorithm is given in Alg. 1.
Input: back-projected depth points of the object $\{q_{m}\}_{m=1}^{M}$; background-masked depth image $D$; initial guesses $\hat{t}$, $\hat{s}$, and $\hat{\mathbf{c}}$; a set of G rotation guesses $\{\hat{R}_{g}\}_{g=1}^{G}$; an ASM with bases $\{b_{k}\}_{k=1}^{K}$; a set of symmetry operations $\{T_{\psi}\}_{\psi=1}^{\Psi}$
Output: estimates of object pose and shape code $\hat{R}$, $\hat{t}$, $\hat{s}$, $\hat{\mathbf{c}}$
$\{\hat{t}_{g},\hat{s}_{g},\hat{\mathbf{c}}_{g}\}\leftarrow\{\hat{t},\hat{s},\hat{\mathbf{c}}\}~{}\forall~{}g$;
for $\text{iter}=1,2,\ldots,\text{maxIterations}$  do
       if down-select hypotheses then
             for $g=1,\ldots,G$ /* execute in parallel */
             do
                   $\mathcal{S}_{tot,g}=\mathcal{S}_{tot}(\hat{R}_{g},\hat{t}_{g},\hat{s}_{g},\hat{\mathbf{c}}_{g};\{q_{m}\}_{m=1}^{M},D,\{T_{\psi}\}_{\psi=1}^{\Psi})$;
             end for
            $\text{sort}(\{\hat{R}_{g},\hat{t}_{g},\hat{s}_{g},\hat{\mathbf{c}}_{g}\}_{g=1}^{G})~{}\text{by}~{}\{\mathcal{S}_{tot,g}\}_{g=1}^{G}$;
             pick top $\text{pick top hypotheses}~{}\{\hat{R}_{g},\hat{t}_{g},\hat{s}_{g},\hat{\mathbf{c}}_{g}\}_{g=1}^{G_{new}}$;
             $G\leftarrow G_{new}$;
       end if
      for $g=1,\ldots,G$ /* execute in parallel */
       do
[…]
             if stop shape estimation then
                   continue;
             end if
            for $\text{step}=1,2,\ldots,\text{maxSteps}$ do
                   find the $\text{find the }Q\text{ closest points }\mathcal{N}_{Q}(q_{m})~{}\forall~{}q_{m}$;
[…]
                   /* (8) instead of (7) in $\mathcal{L}_{sc}$/
             end for
       end for
end for
return $\{\hat{R}_{g=1},\hat{t}_{g=1},\hat{s}_{g=1},\hat{\mathbf{c}}_{g=1}\}$
Algorithm 1 ShapeICP
IV Experiments
IV-A Experiment Setup
Implementation. We implement our ShapeICP method with PyTorch [41] and PyTorch3D [33]. For the segmented depth image of an object, we perform statistical outlier removal on the back-projected point cloud as in [6, 42]. Moreover, we discard object detections that have fewer than a threshold number of points. All the hyperparameter values are summarized in the Appendix.
Datasets. ShapeNetCore [35] is used to compute the ASM for the six categories in the NOCS REAL data [21]. For each category, we exclude weird-looking shapes (e.g. cups with saucers), which is similarly done in [6]. As our model does not train on any pose-annotated data, synthetic benchmarks are irrelevant. We thus evaluate our method on the NOCS REAL test set. For the shape classification network, we use the NOCS REAL train set for training.
Baselines. We benchmark our ShapeICP method against several interesting baselines. NOCS [21] is the seminal work in this research area. Neural Analysis-by-Synthesis [9] is an optimization-based method that is not learning-based and does not rely on image sequences or strong application-specific constraints (see Section II-A for more details). Metric Scale [11] is a rare case of adopting a mesh-based shape representation. ASM-Net [6] also uses ASM, but it is point-based. Unlike our method, its inference is mainly performed by neural network regression. Shape Prior [1] also estimates the shape by deforming a prior, but the method is learning-based. CenterSnap [22] is a one-shot end-to-end learning-based solution, which is the opposite of our work.
Evaluation protocol and metrics. We use the original public NOCS evaluation code for the pose evaluation, as was done in most prior work. A small bug in the NOCS code was later reported by [8]. We provide the updated results with the bug fixed in the Appendix. We use the Chamfer distance as the metric for the shape evaluation and follow the evaluation code from [1]. All the pose accuracy metrics in Tables I and III are in terms of the percentage of the results that fall under the specified thresholds. IoU is the 3D overlap between the estimated 3D bounding box and the ground-truth 3D bounding box of the object. Most neural methods directly predict three scales along the three axes ($x,y,z$) for more accurate 3D bounding box prediction. To build the 3D bounding box, our method multiplies the estimated scale by the aspect ratios along the three axes of the estimated shape, whose diagonal is normalized to unity (the final estimated scale absorbs the normalization factor).
IV-B Results
Table I shows the benchmarking results on NOCS REAL. Our method achieves overall performance that surpasses or is comparable to many data-driven approaches. As an optimization-based approach, our method can reach very high accuracy once it finds the convergence basin around the ground-truth, which is evidenced by its dominating performance in the 5°5cm and 5°10cm metrics. The inability of our method to estimate three scales along the three axes is potentially the cause of the low IoU50 metric.
TABLE I: Benchmarking results on NOCS REAL [21]. Baseline results are from sources [23, 6]. S.D. stands for shape-annotated data required for training and P.D. stands for pose-annotated data required for training. S.C. stands for the shape classification network. The results are roughly in the order of overall performance. The best result in each metric is in bold.
Method P.D. S.D. IoU25 IoU50 IoU75 5°5cm 5°10cm 10°5cm 10°10cm
NOCS [21] 84.8 78.0 30.1 10.0 9.8 25.2 25.8
Synthesis [9] 0.9 1.4 2.4 5.5
Metric Scale [11] 81.6 68.1 5.3 5.5 24.7 26.5
ASM-Net w/o ICP [6] 64.4 31.7 12.4 37.7
ShapeICP(Ours) w/o S.C. 82.6 58.3 37.0 32.4 33.2 42.2 43.4
ASM-Net w/ ICP [6] 68.3 37.4 25.6 43.7
ShapeICP(Ours) w/ S.C. 82.1 58.0 42.2 36.5 37.1 50.6 51.4
ShapePrior [1] 81.2 77.3 53.2 21.4 21.4 54.1 54.1
CenterSnap [22] 83.5 80.2 27.2 29.2 58.8 64.4
TABLE II: Shape estimation results on NOCS REAL [21] evaluated with Chamfer distance ($10^{-3}$). Baseline results are from sources [23, 6]. See the Appendix for per-class results.
Method P.D. S.D. Mean
Categorical Prior [1] 4.4
ShapeICP(Ours) w/o S.C. 3.8
ShapePrior [1] 3.2
ShapeICP(Ours) w/ S.C. 2.6
CenterSnap [22] 1.5
ASM-Net [6] 0.29
We emphasize the comparison between our method and ASM-Net [6] since we both adopt ASM. Despite not using any pose-annotated data, our method surpasses ASM-Net, which fully leverages all the annotated data, by a significant margin. Our no-learning variation (w/o S.C.) is even stronger than their ICP-assisted version in the 5°5cm metric and closely comparable in the IoU75 and 10°5cm metrics. Moreover, we attain a dominant win against the previous no-learning method, Neural Sythesis-by-Analysis [9], and the previous mesh-based method [11].
In Table II, our method provides moderate shape estimation accuracy, ranking in the middle among several learning-based methods. ASM-Net achieves better shape accuracy than ours. The point-based ASM used by ASM-Net is more flexible (due to the undefined connectivity between points) than our mesh-based ASM though it is less expressive in the geometry (due to the loss of surface). The main limitation for our mesh-based ASM comes from the template deformation step which restricts the mesh topology to the template topology, bringing in another source of error other than the latent shape reconstruction error. However, this limitation can be alleviated by manually engineering category-specific templates instead of a spherical template throughout. Additionally, our shape initialization network performs classification instead of regression. It has the advantage of being failure-safe since even if it makes a wrong prediction, the prediction is always a valid shape in the category. However, this merit comes at the expense of losing the ability to interpolate or extrapolate, resulting in lower accuracy.
Figure 4: Visualization of example NOCS REAL [21] frames. From left to right are the color image, the color image overlaid by the depth projection of the estimated objects, the back-projected measured depth image, and the estimated objects from a novel viewpoint.
In Fig. 4, we visualize some example results. The estimated objects in general fit the depth measurements very well, showing the effectiveness of ShapeICP in minimizing the objective function (7). The errors mostly stem from the inherent ambiguity in the problem. For example, a can is easily fit upside down. In addition, our method notably tends to make mistakes on mugs and cameras. The orientation of these objects mostly depends on local features such as the mug handle. Therefore, the global minimum for the orientation has a very small convergence basin, which is extremely difficult to find under the influence of noise and partial view. In general, our algorithm has a tendency to fit the bulk of the object well but ignore the local details.
The current focus of our algorithm is on accuracy, but a discussion of runtime is in the Appendix.
IV-C Ablation
TABLE III: Ablation results on NOCS REAL [21].
Ablated Module IoU75 5°2cm 5°5cm 10°2cm 10°5cm
EM -1.5 -2.1 -4.0 -0.4 +0.2
Symm. Check -0.3 -0.6 -0.5 -0.4 -0.4
Rendering -0.7 -2.5 -2.2 -0.5 +0.2
Multi-Hypotheses -9.9 -4.9 -6.2 -7.8 -9.3
We ablate the modules for local minimum handling introduced in Section III-D in Table III. The symmetry check provides uniform improvements across the metrics. The EM treatment and the depth rendering check lead to more notable improvements in the higher accuracy metrics, such as 5°5cm and 5°2cm. The proposed multi-hypothesis estimation significantly outperforms single-hypothesis estimation that uses $\mathcal{S}_{r}+\mathcal{S}_{\sigma}$ to select the best hypothesis at the beginning of the iterations. More analysis is in the Appendix.
IV-D ShapeICP versus Learning-based Methods
As discussed in Section IV-B, due to the inherent ambiguity in the problem, the estimated pose and shape may fit the depth measurements very well (i.e. achieving a low objective value or even the global minimum) but are different from the ground-truth. This is a limitation of the objective function. Another limitation of our ShapeICP solver is its inability to find a very small convergence basin. These limitations are where learning-based methods can help, since they essentially solve a mapping problem from inputs to outputs instead of an optimization problem. However, our optimization-based approach shines when it finds the convergence basin, it can achieve very high accuracy. Furthermore, our method can be quickly deployed in new environments without curating an object pose dataset for training.
V Conclusion
We proposed the ShapeICP algorithm for category-level object pose and shape estimation from a single depth image. Our method implements a novel mesh-based active shape model (ASM) to represent categorical object shapes. Remarkably, the method does not require pose-annotated data, unlike the vast majority of the methods in the field. However, the performance of ShapeICP surpasses or is on par with many learning-based methods that rely on pose-annotated data. Future directions include how to better disambiguate the problem, potentially by exploring RGB-image-based constraints, and how to better initialize the algorithm closer to the convergence basin of a global minimum.